OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $9 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $132 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 132 = 1/2x + 2(x - 9) What is the regular price of a child's ticket?

OpenStudy (campbell_st):

so if you simplify the equation 132 = 1/2x + 2x - 18 or 132 = 5/2 x - 18 add 18 to both sides 150 = 5/2 x solve for x this will be the price of an adult ticket... then subtract 9 dollars for the childs ticket price...

OpenStudy (anonymous):

so i divide 5/2 from 150

OpenStudy (campbell_st):

start simple... multiply both sides of the equation by 2 that gets rid of the fraction... then divide both sides by 5... that will give the value of the adult ticket

OpenStudy (anonymous):

im so confused

OpenStudy (campbell_st):

ok... so \[150 = \frac{5x}{2}\] multoply both sides by 2.. what do you get...?

OpenStudy (anonymous):

300=10

OpenStudy (campbell_st):

lets go with 300 = 5x does that make sense...?

OpenStudy (anonymous):

kinda

OpenStudy (campbell_st):

if it does make sense now divide both sides by 5 what do you get..?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

final answer would be 51

OpenStudy (campbell_st):

well it would be $51... make sure you include units...

OpenStudy (anonymous):

After a discount of 20%, the price of fitness equipment, including tax, was $73.53. The following equation can be used to find x, the price of the equipment without the discount. x - 0.2x + 0.07(x - 0.2x) = 73.53 What is the price of the fitness equipment without the discount?

OpenStudy (campbell_st):

well collect like terms x - 0.2x + 0.07x - 0.014x = 73.53 or (1 - 0.2 + 0.07 - 0.014)x = 73.53 simplify and then solve

OpenStudy (anonymous):

what

OpenStudy (campbell_st):

ok... get a calculator and add the information in the brackets then divide 73.53 by your answer...

OpenStudy (anonymous):

858.9953271

OpenStudy (anonymous):

right

OpenStudy (anonymous):

Damien works part-time from home as a data entry professional. He enters data in forms and is paid based on the entry accuracy rate. The table below shows his compensation. Based on his regular performance, the equation below can be used to determine C, the total amount Damien earns a month entering data, where x represents the total number of forms he completes that month. C = 14(0.6x) + 12(0.3x) + 0(0.1x) What is the minimum number of forms he must complete in a month to earn a total of at least $1,080? 90 91 1068 1092